# Diffusion-based RGB-D Semantic Segmentation with Deformable Attention Transformer

### v1

@@ Abstract
Vision-based perception and reasoning is essential for scene understanding in any autonomous system. RGB and depth images are commonly used to capture both the semantic and geometric features of the environment. Developing methods to reliably interpret this data is critical for real-world applications, where noisy measurements are often unavoidable. In this work, we introduce a diffusion-based framework to address the RGB-D semantic segmentation problem. Additionally, we demonstrate that utilizing a Deformable Attention Transformer as the encoder to extract features from depth images effectively captures the characteristics of invalid regions in depth measurements. Our generative framework shows a greater capacity to model the underlying distribution of RGB-D images, achieving robust performance in challenging scenarios with significantly less training time compared to discriminative methods. Experimental results indicate that our approach achieves State-of-the-Art performance on both the NYUv2 and SUN-RGBD datasets in general and especially in the most challenging of their image data. Our project page will be available at https://diffusionmms.github.io/
@@ I Introduction
Furthermore, to the best of the authors’ knowledge, all methods on RGB-D semantic segmentation only employ the discriminative-based paradigm [5, 9, 1, 10]. Recently, generative-based methods such as diffusion models have been shown to achieve impressive results on several vision tasks. Although diffusion was originally designed for the image generation problem, many efforts have been made to apply it to RGB semantic segmentation [13, 14, 15].
Motivated by its potential, in this paper we introduce a general, simple, yet effective diffusion framework for high-performance RGB-D semantic segmentation, overcoming several of the identified challenges in the domain. In summary, our contributions are as follows: First, we demonstrate that using a deformable transformer as an image encoder in a discriminative-based architecture can alleviate the problem caused by invalid pixels in depth images. Second, we demonstrate that the use of diffusion can achieve improved results - compared to the State-of-the-Art (SOTA) - combined with reduced training time. Third, experimental results demonstrate that our method achieves SOTA on both the NYUv2 and SUNRGBD datasets and several challenging setups.
@@ II-A RGB-D Semantic Segmentation
Second, the exchange-base fusion strategy involves the exchange of information between modalities through shared representations. The idea is to dynamically refine features from each modality based on insights gained from the other. TokenFusion [10] employs the prune-then-substitute scheme to replace uninformative tokens with more valuable ones from the other modality. This work is studied further by Jia et al [1], where it is suggested that the risk of all tokens engaging in unnecessary exchange can lead to significant information loss, and proposing an exchange-based strategy based on cross-modal transformers instead.
@@ II-B Diffusion in Image Segmentation
All current literature in the field of RGB-D semantic segmentation follows the discriminative paradigm which broadly represents the community standard for semantic segmentation. However, recently generative models have taken the community by storm with their remarkable performance in the image generation task. Several studies suggest that generative models can achieve promising results on segmentation tasks [14, 15, 13]. Baranchuk et al [13] illustrated that feature representations learned from pre-trained diffusion models can be fine-tuned for the segmentation task and achieve SOTA results with limited labels. In [15], Chen et al demonstrated a general framework based on diffusion models for panoptic segmentation on images and videos. In [14], Ji et al proposed a diffusion-based architecture that is suitable for several visual dense prediction tasks including semantic segmentation, depth estimation, and BEV map segmentation. These results are based on RGB image data. Their promising results in combination with the benefits of multi-modality motivated the incorporation of diffusion in the proposed architecture for RGB-D semantic segmentation.
@@ III-A Preliminaries
Figure 2: The architecture of our RGBD semantic mask generation framework. A Deformable Attention Transformer is used as the hierarchical encoder to extract features from RGB and depth images. Multi-scale features from both branches are then processed using fusion modules followed by a Feature Pyramid Network to create a conditioning signal $x$ that matches the shape of the noisy segmentation ground-truth. A deformable attention mask encoder is trained to gradually denoise the concatenated signal to generate the segmentation mask.
@@ III-B Proposed architecture
We start by considering the method described in [9] as the baseline. In [9], Zhang et al proposed a sophisticated fusion module to facilitate interactions between RGB and depth images. However, in order to achieve good results, it relies on the three-channel HHA encoding [12] of the depth images. To understand the underlying challenges of the method, we retrained their models on raw depth images with a high number of invalid pixels which can be up to over 80% the total number of pixels in a depth image and found that the training loss contains multiple spikes (see Figure 3) suggesting training instability and possible overfiting to the noise in the depth images. We argue that instead utilizing the Deformable Attention Transformer (DAT) as the encoder – with its characteristic of having the adaptive spatial aggregation conditioned by input and task information [29]– is well suited to tackle the challenge posed by the invalid pixels in depth images. Furthermore, a diffusion model, as a denoising process, is advantageous when learning the underlying distribution of depth images given their nature of having uncertain measurements.
Figure 3: Loss comparison between training the original CMX [9] model and when using DAT++-S as encoder on raw depth images.
@@ III-B2 Mask Decoder
The mask decoder $f_{\epsilon}$ takes the noisy mask $y_{t}$ and the conditioning signal via concatenation as input. It is then trained to reconstruct the segmentation groundtruth with the original size using the standard cross entropy loss for the semantic segmentation task. Following [14], we simply stack six layers of deformable attention blocks [30] with time embedding as the mask decoder.
@@ III-B3 Training
Label encoding. Diffusion models were originally designed to work with continuous data and Gaussian noise. Several studies have investigated ways to apply diffusion to tasks with discrete labels [31, 15, 14] such as semantic segmentation. Inspired by the work of Ji et al [14], we use the class embedding approach in which a learnable embedding layer is used to project discrete labels into a high-dimensional space, normalized by a sigmoid function. The encoded labels are scaled to the range of $[-s,s]$ as shown in Algorithm 1.
@@ III-B4 Inference
The inference process is outlined in Algorithm 2. Given paired RGB and depth images as the conditional inputs, the model begins with a random noise map generated from a Gaussian distribution and progressively improves the prediction. To minimize the number of iterative steps, we choose the DDIM update rule [28, 14] for the sampling process. The key hyperparameter in this update rule is the time difference td which determines how far apart consecutive timesteps are chosen during the reverse diffusion process. Larger time gaps between steps allow for faster sampling at the potential cost of sample quality. Our experiments show that td${}=1$ and the number of steps of 3 work best for our method. The details of the DDIM update rule is presented in Algorithm 3.
@@
mask_next =$\sqrt{\alpha_{next}}$* mask_enc +$\sqrt{1-\alpha_{now}}$*eps
@@ IV-A2 Training details and hyperparameters
For all experiments, we train our models using the AdamW optimizer [36], with an initial learning rate of $6\times 10^{-5}$ and a weight decay of 0.01. The learning rate is regulated using the warmup polynomial decay scheduler. As seen in many previous works [9, 20, 5], the discriminative paradigm needs at least 300 to 500 epochs to obtain good results on both NYUv2 and SUN-RGBD datasets. We found that our diffusion-based model can converge in significantly less amount of time. We train the NYUv2 and SUN-RGBD datasets for 100 and 50 epochs, respectively. For data augmentation, we perform resize, random crop, and random flip operations on both RGB and depth images.
IV-A3 Ablation on architectures
Figure 4 depicts reference sampling points learned by the offset network in the Deformable Attention Transformer encoder. For each attention module, reference points are initially created as uniform grids that remain consistent across the input data. An offset network then processes all query features to generate corresponding offsets for these reference points. They guide the candidate keys/values toward important regions, enhancing the original self-attention module with greater flexibility and efficiency to capture more important features. It can be seen in Figure 4 that the learned reference points are shifted towards salient regions in the depth image, including the large invalid region covering the microwave in the RGB image. By avoiding the discontinuous transition between valid and invalid regions, the model can perceive the invalid regions as a whole object instead of overfitting to noisy measurements in the depth image. This insight might hold true for invalid regions caused by other reflective surfaces such as television screens or windows suggesting that DAT, with its data-dependent characteristic, can effectively exploit depth features even under severe uncertainty measurements.
Figure 4: Reference sampling points
@@ IV-A3 Ablation on architectures
The superiority of DAT on RGB-D semantic segmentation over other non data-dependent encoders such as MixTransformer [37] is quantitatively shown in Table I. We take the result reported in [9] as the baseline. By simply changing the encoder of the model to DAT++-S [24], we achieved 3.3% meanIoU boost in the NYUv2 dataset as seen in Table I. However, when evaluating a larger dataset like SUN-RGBD using the discriminative model, we did not achieve any performance boost. By using the diffusion model as the alternating paradigm, we achieved better results compared to the discriminative model on both datasets as seen in Table I while spending significantly less time/epochs to train.
@@
TABLE I: Performance comparison between DAT++ and MiT-B5 using different paradigms on NYUv2 and SUN-RGBD dataset
Methods NYUv2 SUN-RGBD
CMX (MiT-B5) 56.9 52.4
Discriminative-based (DAT++-S[24]) 60.2 52.4
Diffusion-based (DAT++-S[24]) 61.2 53.7
IV-A4 Evaluation results on challenging datasets
@@
We evaluate our diffusion based framework using 3 encoder variants namely DAT++-T, DAT++-S, DAT++-B [24]. They achieve SOTA results compared to current literature in the field as shown in Table II. On the NYUv2 dataset, our method with the DAT++-S encoder achieves the best meanIoU of 61.2% which is the number one on the dataset leaderboard at the time of writing. Interestingly, increasing the model size to DAT++-B leads to 0.4% performance drop. This might be the overfitting problem when using a large model on a small dataset like NYUv2 dataset. On the SUN-RGBD dataset, we achieved the best result of 53.9% with DAT++-B encoder. We lose only to GeminiFusion with Swin-Large encoder (197M parameters) which is much larger than DAT++-B encoder (94M parameters). Therefore, our method remains the best method performing on the SUN-RGBD dataset compared with other similar size models. Furthermore, when we evaluate on the SUN-RGBD dataset using the model trained on NYUv2 dataset and vice versa, we observe only a modest performance drop for all variants of our model, which suggests a good generalization capability of our method.
@@
TABLE II: Comparison with other State-of-the-art methods on the NYUv2 and SUN-RGBD datasets. Red numbers indicate the results when evaluating SUN-RGBD dataset using the model trained on NYUv2 and vice versa. The scale value $s=0.01$
Methods NYUv2 SUN-RGBD
CMX (MiT-B4)[9] 56.3 52.1
CMX (MiT-B5)[9] 56.9 52.4
DFormer-S[5] 53.6 50.0
DFormer-B[5] 56.9 51.2
DFormer-L[5] 57.2 52.5
TokenFusion 54.2 52.4
GeminiFusion (MiT-B3)[1] 56.8 52.7
GeminiFusion (MiT-B5)[1] 57.7 53.3
GeminiFusion (Swin-L)[1] 60.9 54.6
Ours (DAT++-T) 59.7 (59.7) 52.6 (50.6)
Ours (DAT++-S) 61.2 (61.1) 53.7 (51.2)
Ours (DAT++-B) 60.8 (60.4) 53.9 (51.4)
@@
Figure 5: Our segmentation results show that our method has better understanding of invalid depth regions than other methods in both normal (upper row) and low-light conditions (lower row)
@@
TABLE III: Comparison with other methods on challenging subsets. The red numbers indicate the relative performance drop compared with evaluation results on original datasets. DAT++-S is used as the encoder
@@
Ours (s=0.01) 61.2 58.8 (2.4) 59.3 (1.9) 56.6 (4.6)
Ours (s=0.03) 61.5 58.9 (2.6) 58.7 (2.8) 56.6 (4.9)
@@
Ours (s=0.01) 53.7 52.2 (1.5) 48.2 (5.5) 49.0 (4.7)
Ours (s=0.03) 53.8 52.3 (1.5) 47.8 (6.0) 49.0 (4.8)
To further demonstrate the capacity of our method to exploit depth features under heavy uncertainty, we conduct further evaluations on three focused of sub-datasets, namely
  • Low light: For each paired RGB-D frame in the original datasets, we deliberately decrease the intensity of the RGB images using gamma correction operation. The intensity of a normalized image $I$ is adjusted based on the formula $I_{dark}=I^{\gamma},\gamma=2$. We call this the ‘low-light’ subset of the dataset.
  • Small objects: We ignore some labels in the ground truth which are less affected by noisy depth data and only evaluate the meanIoU metric of the rest. For NYUv2, we ignore the [“wall”, “floor”, “ceiling”, “otherstructure”, “otherfurniture”, “otherprop”] labels. For SUN-RGBD, we ignore the [“wall”, “floor”, “ceiling”] labels. We call this the ‘small’ objects dataset.
Table III shows that our method outperforms others across all three challenging datasets, with less performance drop. This proves its effectiveness in modeling RGB-D images, even with a high percentage of invalid depth pixels. Qualitative results are shown in Figure 5
IV-A5 Ablation study
We conduct ablation experiments to find the best value for several key hyperparameters in diffusion models. Table IV shows that using cosine schedule achieves slightly better results than linear schedule.
@@ IV-A5 Ablation study
Table V shows that 0.03 and 0.001 are the best scale value for NYUv2 and SUN-RGBD dataset respectively. It is noted that $s=0.01$ and $s=0.03$ work well for both datasets, whereas $s=0.001$ improves the performance on SUN-RGBD by a small margin but disproportionally worsens the results on the NYUv2 dataset.

### v3

@@ Abstract
Vision-based perception is of great importance for scene understanding in autonomous systems. RGB-D images are commonly used to capture both semantic and geometric features, but reliable interpretation is challenging due to unavoidable noise in real-world data. In this work, we introduce a diffusion-based framework to address the RGB-D semantic segmentation problem. Additionally, we demonstrate that utilizing a Deformable Attention Transformer as the encoder to extract features from depth images effectively captures the characteristics of invalid regions in depth measurements. Our generative framework shows a greater capacity to model the underlying distribution of RGB-D images, achieving robust performance in challenging scenarios with significantly less training time compared to discriminative methods. Experimental results indicate that our approach achieves State-of-the-Art performance on both the NYUv2 and SUN-RGBD datasets in general and especially in the most challenging of their image data. To demonstrate the practicality of our method, a real-world experiment is conducted to inspect an office and generate its 3D semantic map. Our project page will be available at https://diffusionmms.github.io/
@@ I Introduction
Furthermore, to the best of the authors’ knowledge, all methods on RGB-D semantic segmentation only employ the discriminative paradigm [5, 9, 1, 10]. Recently, generative-based methods such as diffusion models have been shown to achieve impressive results on several vision tasks. Although diffusion was originally designed for the image generation problem, many efforts have been made to apply it to RGB semantic segmentation [13, 14, 15].
Motivated by its potential, we propose a simple yet effective diffusion framework for high-performance RGB-D semantic segmentation that addresses key domain challenges. In summary, our contributions are as follows: First, we demonstrate that using a deformable transformer as an image encoder in a discriminative-based architecture can alleviate the problem caused by invalid pixels in depth images. Second, we demonstrate that the use of diffusion can achieve improved results - compared to the State-of-the-Art (SOTA) - combined with reduced training time. Third, experimental results demonstrate that our method achieves SOTA on both the NYUv2 and SUNRGBD datasets and several challenging setups. We also validate our approach with a real-world drone experiment reconstructing a 3D semantic map of an office.
@@ II-A RGB-D Semantic Segmentation
Second, the exchange-based fusion strategy involves the exchange of information between modalities through shared representations. The idea is to dynamically refine features from each modality based on insights gained from the other. TokenFusion [10] employs the prune-then-substitute scheme to replace uninformative tokens with more valuable ones from the other modality. This work is studied further by Jia et al [1], where it is suggested that the risk of all tokens engaging in unnecessary exchange can lead to significant information loss, and proposing an exchange-based strategy based on cross-modal transformers instead.
@@ II-B Diffusion in Image Segmentation
All current literature in the field of RGB-D semantic segmentation follows the discriminative paradigm which broadly represents the community standard for semantic segmentation. However, generative models have recently taken the community by storm with their remarkable performance in the image generation task. Several studies suggest that generative models can achieve promising results on segmentation tasks [14, 15, 13]. Baranchuk et al [13] illustrated that feature representations learned from pre-trained diffusion models can be fine-tuned for the segmentation task and achieve SOTA results with limited labels. In [15], Chen et al demonstrated a general framework based on diffusion models for panoptic segmentation on images and videos. In [14], Ji et al proposed a diffusion-based architecture that is suitable for several visual dense prediction tasks, including semantic segmentation, depth estimation, and BEV map segmentation. These results are based on RGB image data. Their promising results in combination with the benefits of multi-modality motivated the incorporation of diffusion in the proposed architecture for RGB-D semantic segmentation.
@@ III-A Preliminaries
Figure 2: The architecture of our RGBD semantic mask generation framework. A Deformable Attention Transformer is used as the hierarchical encoder to extract features from RGB and depth images. Multi-scale features from both branches are then processed using fusion modules followed by a Feature Pyramid Network to create a conditioning signal $x$ that matches the shape of the noisy segmentation ground truth. A deformable attention mask encoder is trained to gradually denoise the concatenated signal to generate the segmentation mask.
@@ III-B Proposed architecture
We start by considering the method described in [9] as the baseline. In [9], Zhang et al proposed a sophisticated fusion module to facilitate interactions between RGB and depth images. However, in order to achieve good results, it relies on the three-channel HHA encoding [12] of the depth images. To understand the underlying challenges of the method, we retrained their models on raw depth images with a high number of invalid pixels which can be up to over 80% the total number of pixels in a depth image and found that the training loss exhibits multiple spikes (see Figure 3) suggesting training instability and possible overfitting to the noise in the depth images. We argue that instead utilizing the Deformable Attention Transformer (DAT) as the encoder – with its characteristic of having the adaptive spatial aggregation conditioned by input and task information [29]– is well suited to tackle the challenge posed by the invalid pixels in depth images. Furthermore, a diffusion model, as a denoising process, is advantageous when learning the underlying distribution of depth images given their nature of having uncertain measurements.
Figure 3: Loss comparison between training the original CMX [9] model and when using DAT++-S as the encoder for raw depth images.
@@ III-B2 Mask Decoder
The mask decoder $f_{\epsilon}$ takes the noisy mask $y_{t}$ and the conditioning signal via concatenation as input. It is then trained to reconstruct the segmentation ground truth with the original size using the standard cross-entropy loss for the semantic segmentation task. Following [14], we form the mask decoder by stacking six layers of deformable attention blocks [30] with time embedding.
@@ III-B3 Training
Label encoding. Diffusion models were originally designed to work with continuous data and Gaussian noise. Several studies have investigated ways to apply diffusion to tasks with discrete labels [31, 15, 14] such as semantic segmentation. Inspired by the work of Ji et al [14], we use the class embedding approach in which a learnable embedding layer is used to project discrete labels into a high-dimensional space, normalized by a sigmoid function. The encoded labels are scaled to the range of [s,s][-s,s] as shown in Algorithm 1.
@@ III-B4 Inference
The inference process is outlined in Algorithm 2. Given paired RGB and depth images as the conditional inputs, the model begins with a random noise map generated from a Gaussian distribution and progressively improves the prediction. To minimize the number of iterative steps, we choose the DDIM update rule [28, 14] for the sampling process. The key hyperparameter in this update rule is the time difference td which determines how far apart consecutive timesteps are chosen during the reverse diffusion process. Larger time gaps between steps allow for faster sampling at the potential cost of sample quality. Our experiments show that td=1~=1 works best for our method. The details of the DDIM update rule are presented in Algorithm 3.
@@
mask_next $\sqrt{\alpha_{next}}*$ mask_enc $\sqrt{1-\alpha_{now}}*$eps
@@ IV-A1 Datasets and Metrics
Figure 4: GradCAM (Gradient-weighted Class Activation Map) heatmaps visualizing class activations from the final layer of the backbone used in our method and in CMX [9]. For each method, we present (from left to right) the predicted semantic segmentation mask, the binary mask for the selected class, and the corresponding GradCAM heatmap. The top row displays results for the class ’window’. The bottom row shows results for the class ’table’.
@@ IV-A2 Training details and hyperparameters
For all experiments, we train our models using the AdamW optimizer [36], with an initial learning rate of $6\times 10^{-5}$ and a weight decay of 0.01. The learning rate is regulated using the warmup polynomial decay scheduler. As seen in many previous works  [9, 5, 20], the discriminative paradigm needs at least 300 to 500 epochs to obtain good results on both NYUv2 and SUN-RGBD datasets. We found that our diffusion-based model can converge in a significantly less amount of time. We train the NYUv2 and SUN-RGBD datasets for 100 and 50 epochs, respectively. For data augmentation, we perform resize, random crop, and random flip operations on both RGB and depth images.
IV-B Ablation study
IV-B1 Comparison with the baseline CMX [9]
We adopt the fusion mechanism from CMX [9] as the foundation of our architecture, but introduce two key enhancements to improve performance on depth images with invalid regions. First, we replace encoders used in CMX [9] with a Deformable Attention Transformer encoder that adaptively attends to spatially relevant regions in both RGB and depth images. Second, we shift the training paradigm for semantic segmentation from a discriminative approach to a generative one.
@@ IV-B1 Comparison with the baseline CMX [9]
To visualize the effectiveness of the Deformable Attention Transformer encoder in obtaining more robust features of depth images, we use the Gradient-weighted Class Activation Mapping (GradCAM) [37] technique to highlight the regions in an input image that are most influential for a model’s decision. As shown in Figure 4, the Deformable Attention Transformer encoder can handle invalid regions caused by reflective surfaces such as windows or glass better than encoders used in CMX [9]. The superiority of DAT on RGB-D semantic segmentation over other non data-dependent encoders such as MixTransformer [38] is quantitatively shown in Table I. We take the result reported in [9] as the baseline. By simply changing the encoder of the model to DAT++ [24], we achieved a significant boost in terms of the meanIoU metric in the NYUv2 dataset, as seen in Table I across all versions of the DAT++ backbone. Note that the number of parameters reported in Table I refers to the size of the backbone network used for feature extraction from each input. Utilizing the tiny version of DAT++ achieves a $1.8\%$ improvement in mean IoU while reducing the number of parameters to one-third. Note that the number of parameters in Table I reflects the size of the encoder used to extract features from each modality. However, when evaluating a larger dataset like SUN-RGBD using the discriminative model, we did not achieve any performance boost. By using the diffusion model as the alternating paradigm, we achieved better results compared to the discriminative model on both datasets, as seen in Table I while spending significantly less time/epochs to train.
Our method inherits the multiple-step inference procedure from the diffusion process. This introduces a tradeoff between model performance and computational cost, depending on the number of sampling steps required for inference. Typically, at least 50 sampling steps are necessary to achieve high-fidelity results in image generation tasks [28]. However, as shown in Figure 5, the performance of the diffusion process for the semantic segmentation task quickly saturates after just a few sampling steps. Even with a single sampling step, our method outperforms the baseline by a significant margin. The reported latency of our models was measured using an NVIDIA GeForce RTX 3090.
@@
TABLE I: Performance comparison between DAT++ and MiT-B5 using different paradigms on NYUv2 and SUN-RGBD datasets. The number of parameters refers to the size of the encoder used to extract features from each modality.
Methods Params(M) NYUv2 SUN-RGBD CMX (MiT-B5) 82 56.9 52.4 Discriminative-based (DAT++-T) 24 58.7 51.7 Discriminative-based (DAT++-S) 53 60.2 52.4 Discriminative-based (DAT++-B) 93 59.9 51.8 Diffusion-based (DAT++-T) 24 59.7 52.6 Diffusion-based (DAT++-S) 53 61.5 53.8 Diffusion-based (DAT++-B) 93 60.8 54.0
IV-B2 Evaluation results on challenging datasets
Figure 5: Performance and computational cost of our method with different numbers of sampling steps used for inference on NYUv2 dataset
@@ IV-B2 Evaluation results on challenging datasets
We evaluate our diffusion based framework using 3 encoder variants, namely DAT++-T, DAT++-S, DAT++-B [24]. They achieve SOTA results compared to current literature in the field, as shown in Table II. On the NYUv2 dataset, our method with the DAT++-S encoder achieves the best meanIoU of 61.2% which is the number one on the dataset leaderboard at the time of writing. Interestingly, increasing the model size to DAT++-B leads to 0.4% performance drop. This might be the overfitting problem when using a large model on a small dataset like the NYUv2 dataset. On the SUN-RGBD dataset, we achieved the best result of 53.9% with the DAT++-B encoder. We lose only to GeminiFusion with the Swin-Large encoder (369.2M parameters), which is much larger than our biggest model with the DAT++-B encoder (256.8M parameters). Therefore, our method remains the best method performing on the SUN-RGBD dataset compared with other similar size. Our method outperforms others more on the small NYUv2 dataset than on the larger SUN-RGBD, showing its strength with limited data. Furthermore, when we evaluate on the SUN-RGBD dataset using the model trained on the NYUv2 dataset and vice versa, we observe only a modest performance drop for all variants of our model, which suggests a good generalization capability of our method.
@@
TABLE II: Comparison with other State-of-the-art methods on the NYUv2 and SUN-RGBD datasets. Red numbers indicate the results when evaluating the SUN-RGBD dataset using the model trained on NYUv2 and vice versa. The scale value $s=0.01$
Methods Backbone Params(M) NYUv2 SUN-RGBD CMX (MiT-B4)[9] MiT-B4 139.9 56.3 52.1 CMX (MiT-B5)[9] MiT-B5 181.1 56.9 52.4 DFormer-S[5] DFormer-S 18.7 53.6 50.0 DFormer-B[5] DFormer-B 29.5 56.9 51.2 DFormer-L[5] DFormer-L 39.0 57.2 52.5 TokenFusion[10] MiT-B3 45.9 54.2 52.4 GeminiFusion [1] MiT-B3 75.8 56.8 52.7 GeminiFusion [1] MiT-B5 137.2 57.7 53.3 GeminiFusion [1] Swin-L 369.2 60.9 54.6 Ours (step 3) DAT++-T 73.2 95.9 (59.7) 52.6 (50.6) Ours (step 3) DAT++-S 172.8 61.2 (61.1) 53.7 (51.2) Ours (step 3) DAT++-B 280.0 60.8 (60.4) 53.9 (51.4)
@@
Figure 6: Our segmentation results show that our method has a better understanding of invalid depth regions than other methods in both normal (upper row) and low-light conditions (lower row)
@@
TABLE III: Comparison with other methods on challenging subsets. The red numbers indicate the relative performance drop compared with evaluation results on the original datasets. DAT++-S is used as the encoder.
@@
Ours (s=0.01) (step 3) 61.2 58.8 (2.4) 59.3 (1.9) 56.6 (4.6)
Ours (s=0.03) (step 3) 61.5 58.9 (2.6) 58.7 (2.8) 56.6 (4.9)
@@
Ours (s=0.01) (step 3) 53.7 52.2 (1.5) 48.2 (5.5) 49.0 (4.7)
Ours (s=0.03) (step 3) 53.8 52.3 (1.5) 47.8 (6.0) 49.0 (4.8)
@@
To further demonstrate the capacity of our method to exploit depth features under heavy uncertainty, we conduct further evaluations on three focused sub-datasets, namely
  • Low light: For each paired RGB-D frame in the original datasets, we deliberately decrease the intensity of the RGB images using a gamma correction operation. The intensity of a normalized image $I$ is adjusted based on the formula $I_{dark}=I^{\gamma},\gamma=2$. We call this the ‘low-light’ subset of the dataset.
  • Small objects: We ignore some labels in the ground truth that are less affected by noisy depth data and only evaluate the meanIoU metric of the rest. For NYUv2, we ignore the [“wall”, “floor”, “ceiling”, “otherstructure”, “otherfurniture”, “otherprop”] labels. For SUN-RGBD, we ignore the [“wall”, “floor”, “ceiling”] labels. We call this the ‘small’ objects dataset.
Table III shows that our method outperforms others across all three challenging datasets, with less performance drop. This proves its effectiveness in modeling RGB-D images, even with a high percentage of invalid depth pixels. Qualitative results are shown in Figure 6. The improved performance in visual degradation cases demonstrates our model’s ability to extract meaningful features directly from raw depth images, avoiding the need for interpolation used by other methods.
IV-B3 Hyperparameter tuning
We conduct ablation experiments to find the best value for several key hyperparameters in diffusion models. Table IV shows that using the cosine schedule achieves slightly better results than the linear schedule. Table V shows that 0.03 and 0.001 are the best scale values for NYUv2 and SUN-RGBD datasets, respectively. It is noted that $s=0.01$ and $s=0.03$ work well for both datasets, whereas $s=0.001$ improves the performance on SUN-RGBD by a small margin but disproportionately worsens the results on the NYUv2 dataset.